40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 15 20:00:23 2020, Koji, Summary, General, HEPA AC cord replacement P_20201015_200732.jpgP_20201015_200752.jpgP_20201015_202615.jpgP_20201015_204234.jpg
    Reply  Thu Apr 1 23:55:33 2021, Koji, Summary, General, HEPA AC cord replacement 
       Reply  Fri Apr 2 15:17:23 2021, gautam, Summary, General, HEPA AC cord replacement 
          Reply  Tue Apr 6 21:17:04 2021, Koji, Summary, General, PSL HEPA investigation 20210406211741_IMG_0554.jpeg20210406211840_IMG_0555.jpeg20210406211850_IMG_0556.jpeg
Message ID: 16002     Entry time: Tue Apr 6 21:17:04 2021     In reply to: 15992
Author: Koji 
Type: Summary 
Category: General 
Subject: PSL HEPA investigation 

- Last week we found both of the PSL HEPA units were not running.

- I replaced the capacitor of the north unit, but it did not solve the issue. (Note: I reverted the cap back later)
- It was found that the fans ran if the variac was removed from the chain.
- But I'm not certain that we can run the fans in this configuration with no attendance considering fire hazard.

@3AM: UPON LEAVING the lab, I turned off the HEPA. The AC cable was not warm, so it's probably OK, but we should wait for the continuous operation until we replace the scorched AC cable.


The capacitor replacement was not successful. So, the voltages on the fan were checked more carefully. The fan has the three switch states (HIGH/OFF/LOW). If there is no load (SW: OFF), the variac out was as expected. When the load was LOW or HIGH, it looked as if the motor is shorted (i.e. no voltage difference between two wires).

I thought the motors may have been shorted. But if the load resistance was measured with the fluke meter, it showed some resistance

- North Unit: SW LOW 4.6Ohm / HIGH 6.0Ohm
- South Unit: SW LOW 6.0Ohm / HIGH 4.6Ohm (I believe the internal connection is incorrect here)

I believed the motors are alive! Then the fans were switched on with the variac removed... they ran. So I set the switch LOW for the north unit and HIGH for the south unit.

Then I inspected the variac:

  • The AC output has some liquid leaking (oil?) (Attachment 1)
  • The AC plug on the variac out has a scorch mark (Attachments 2/3)

So, this scorched AC plug/cable connected directly to the AC right now. I'm not 100% confident about the safety of this configuration.
Also I am not sure what was wrong with the system.

  • Has the variac failed first? Because of the heat? I believe that the HEPA was running @30% most of the time. Maybe the damage was already there at the failure in Nov 2020?
  • Or has the motor stopped at some point and this made the variac failed?
  • Was the cable bad and the heat made the variac failed (then the problem is still there).

So, while I'm in the lab today, I'll keep the HEPA running, but upon my taking off, I'll turn it off. We'll discuss what to do in the meeting tomorrow.

 

Attachment 1: 20210406211741_IMG_0554.jpeg  419 kB  Uploaded Tue Apr 6 22:56:17 2021  | Hide | Hide all
20210406211741_IMG_0554.jpeg
Attachment 2: 20210406211840_IMG_0555.jpeg  167 kB  Uploaded Tue Apr 6 22:56:25 2021  | Hide | Hide all
20210406211840_IMG_0555.jpeg
Attachment 3: 20210406211850_IMG_0556.jpeg  135 kB  Uploaded Tue Apr 6 22:56:28 2021  | Hide | Hide all
20210406211850_IMG_0556.jpeg
ELOG V3.1.3-